# Efficient computation of stationary measures and the Lyapunov Landscape for families random dynamical systems with smooth additive noise

### v1

@@ 8. Fourier Approximation
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left[(1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}+\epsilon\right].\qed$
@@ 9.1. Norm Bounds for the Discretized Operator
We first bound the $L^{1}$ norm of iterates of the discretized operator.
@@
Let $P_{\sigma,K}$ be the discretized annealed operator. Then for all $i\in\mathbb{N}$:
$\|P^{i}_{\sigma,K}\|_{L^{1}\to L^{1}}\leq(1+\Gamma_{\sigma,K})^{i}(2K+1).$
@@
Using $\pi_{K}^{2}=\pi_{K}$ (Lemma 8.3), we write:
$P^{i}_{\sigma,K}=(\pi_{K}N_{\sigma}P)^{i}\pi_{K}.$
By Lemma 8.6, $\|\pi_{K}N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1+\Gamma_{\sigma,K}$, and since $\|P\|_{L^{1}\to L^{1}}=1$, the composition has norm at most $1+\Gamma_{\sigma,K}$. Lastly, since $|\mathcal{F}(g)[j]|\leq\|g\|_{L^{1}}$, we estimate:
$\|\pi_{K}g\|_{L^{1}}\leq\sum_{j=-K}^{K}|\mathcal{F}(g)[j]|\leq(2K+1)\|g\|_{L^{1}}.$
Combining both gives the result. ∎
@@
Let $P_{\sigma}$ be the annealed transfer operator. Then, there exist $K,N\in\mathbb{N}$ such that for all $k>K$:
$\|P^{N}_{\sigma,k}|_{\mathcal{U}_{0}}\|_{L^{2}}\leq 3/4.$
Sketch of proof.
Choose NN such that $\|P_{\sigma}^{N}|_{\mathcal{U}_{0}}\|_{L^{2}}\leq 1/2$ (from Proposition 6.7). Then for any $g\in\mathcal{U}_{0}$:
$\|P_{\sigma,k}^{N}g\|_{L^{2}}$ $\leq\|P_{\sigma}^{N}g\|_{L^{2}}+\|P_{\sigma}^{N}g-P_{\sigma,k}^{N}g\|_{L^{2}}$
$\leq\frac{1}{2}+\text{(discretization error)}.$
The error is expanded using a telescoping sum:
$\sum_{i=0}^{N-1}\|P_{\sigma}^{i}(P_{\sigma}-P_{\sigma,k})P_{\sigma,k}^{N-i-1}g\|_{L^{2}},$
and each term is bounded using Lemma 9.1, Proposition 8.10, and exponential decay of $\Gamma_{\sigma,k}$. The sum is ultimately controlled by:
$\sum_{i=0}^{N-1}C_{i}(1+\Gamma_{\sigma,k})^{N}(2K+1)\Gamma_{\sigma,k},$
which can be made less than $1/4$ for large enough $K$. ∎
@@
For all $i\in\mathbb{N}$, we have:
$\|P_{\sigma,K}^{i}\|_{L^{2}}\leq\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}(1+\Gamma_{\sigma,K})^{i-1}(2K+1)\sqrt{2}.$
Sketch of proof.
The bound follows from combining:
  • Young’s inequality for convolution with $\tau_{*}\rho_{\sigma}$ (see Lemma 6.2),
  • the fact that $\|N_{\sigma}\|_{L^{1}\to L^{2}}=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}$,
  • the estimate $\|\pi_{K}g\|_{L^{1}}\leq(2K+1)\|g\|_{L^{1}}$,
  • and Lemma 9.1 for controlling $L^{1}$ norms of iterates.
@@
Let $f_{\sigma}$ be the fixed point of $P_{\sigma}$, and $f_{\sigma,K}$ the fixed point of $P_{\sigma,K}$. Then for any $\epsilon>0$, there exists $K$ such that:
@@
Apply Theorem 8.10 with the error term
$(1+\Gamma_{\sigma,K}+\|\rho_{\sigma}\|_{L^{2}})\Gamma_{\sigma,K}\|f_{\sigma}\|_{L^{1}},$
multiplied by the norm bounds from Lemma 9.3 and the number of terms NN in Lemma 9.2. Since $\Gamma_{\sigma,K}\to 0$ exponentially as $K\to\infty$, we can ensure the full expression is less than $\epsilon$. ∎
@@
Let $\phi$ be an observable in $L^{2}(m)$; the Birkhoff average
@@
converges almost surely to $\int\phi\,f_{\sigma}dm$, Then for any $\epsilon>0$, there exists $K$ such that:
$\left|\int\phi\,f_{\sigma}dm-\int\phi\,f_{\sigma,K}dm\right|<\epsilon.$
@@
Due to uniqueness of the stationary measure (Proposition 6.8) the convergence of the Birkhoff average follows from Birkhoff Ergodic Theorem 5.12; by the Cauchy-Schwarz inequality we have that
$\left|\int\phi\,f_{\sigma}dm-\int\phi\,f_{\sigma,K}dm\right|\leq\|\phi\|_{L^{2}}||\cdot\|f_{\sigma}-f_{\sigma,K}\|_{L^{2}}$
and the thesis follows from Theorem 9.4. ∎

### v2

@@ 8. Fourier Approximation
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left(\delta+\epsilon\right).$
8.1. Implementation of the discretization for the discretized operator
To discretize the deterministic transfer operator $P$ we use a Fourier–Galerkin projection. Let $e_{k}(x):=e^{k\pi ix}$ and define the matrix coefficients by
$P_{k,\ell}:=\frac{1}{2}\int_{-1}^{1}\overline{e_{k}(x)}\,(Pe_{\ell})(x)\,dx.$
Using the duality relation $\int\varphi\cdot P\psi\,dx=\int(\varphi\circ T)\cdot\psi\,dx$ we obtain
$P_{k,\ell}=\frac{1}{2}\int_{-1}^{1}\overline{e_{k}(T(x))}\,e_{\ell}(x)\,dx.$
For $\alpha>2$, the map $T_{\alpha,\beta}$ is $C^{2}$ on (1,1)(-1,1), hence the function
$f_{k}(x):=\overline{e_{k}(T_{\alpha,\beta}(x))}$
is $C^{2}$ on (1,1)(-1,1). Moreover, while $f_{k}(1)=f_{k}(-1)$ (since $T_{\alpha,\beta}(1)=T_{\alpha,\beta}(-1)$), its derivative need not match at the endpoints; in other words, the 22–periodic extension of $f_{k}$ typically has a jump in the first derivative at $\pm 1$. Consequently, the Fourier coefficients of $f_{k}$ (and hence the entries $P_{k,\ell}$) decay quadratically: for every $\ell\neq 0$,
(16) $|P_{k,\ell}|\;=\;\bigl|\mathcal{F}(f_{k})[\ell]\bigr|\;\leq\;\frac{1}{2(\pi\ell)^{2}}\Bigl(|f_{k}^{\prime}(1)-f_{k}^{\prime}(-1)|+\|f_{k}^{\prime\prime}\|_{L^{1}}\Bigr),$
and in particular $|P_{k,\ell}|\leq C_{k}/|\ell|^{2}$ with an explicit constant $C_{k}$.
For performance we evaluate the Fourier coefficients from point samples using FFTW. Concretely, for a 22–periodic integrand $f$ we sample on an equispaced grid
$x_{m}:=-1+\frac{2m}{N},\qquad m=0,\dots,N-1,$
and compute the discrete Fourier transform of the vector $(f(x_{m}))_{m=0}^{N-1}$ by FFTW. This yields the discrete coefficients
$\mathcal{F}_{N}(f)[\ell]:=\frac{1}{N}\sum_{m=0}^{N-1}f(x_{m})\,e^{-\pi i\ell x_{m}},\qquad|\ell|\leq\frac{N}{2},$
which approximate the true Fourier coefficients
$\mathcal{F}(f)[\ell]=\frac{1}{2}\int_{-1}^{1}f(x)\,e^{-\pi i\ell x}\,dx.$
The discretization error is an aliasing (wrap-around) effect: one has the exact identity
(17) $\mathcal{F}_{N}(f)[\ell]=\sum_{q\in\mathbb{Z}}\mathcal{F}(f)[\ell+qN],$
and hence
(18) $|\mathcal{F}_{N}(f)[\ell]-\mathcal{F}(f)[\ell]|\leq\sum_{q\neq 0}|\mathcal{F}(f)[\ell+qN]|.$
In our application $f=f_{k}:=\overline{e_{k}\circ T}$ is $C^{2}$ on (1,1)(-1,1) and its 22–periodic extension typically has a jump in the first derivative at $\pm 1$; consequently $|\mathcal{F}(f_{k})[n]|\leq C_{k}/|n|^{2}$ for $n\neq 0$ with explicit $C_{k}$ (see (16)). Combining this decay with (18) gives the oversampling bound
(19) $|\mathcal{F}_{N}(f_{k})[\ell]-\mathcal{F}(f_{k})[\ell]|\leq\frac{\pi^{2}}{3}\,\frac{C_{k}}{N^{2}},\qquad|\ell|\leq\frac{N}{2}.$
We therefore choose an oversampling factor $s>1$ and take $N=2^{n}$ (so that we are in the radix-$2$ setting of Lemma 8.18 and of the roundoff analysis in [11]), with $N\geq s(2K+1)$, compute $\mathcal{F}_{N}(f_{k})$ by FFT, and retain the modes $|\ell|\leq K$.
We refer to [11] for a detailed modern analysis of rounding errors in radix-$2$ Cooley–Tukey FFTs. As already emphasized by [31], the comparatively low operation count of the FFT (only $O(N\log N)$ floating-point operations) yields normwise error bounds with small, explicitly controlled constants.
Lemma 8.18.
Let $x\in\mathbb{C}^{N}$ and let
$\mathcal{F}_{N}:=N^{-1/2}\bigl(\omega_{N}^{jk}\bigr)_{j,k=0}^{N-1},\qquad\omega_{N}:=e^{-2\pi i/N},$
be the unitary Fourier matrix. Denote by $y:=\mathcal{F}_{N}x$ the exact transform and by $\widehat{y}$ the result returned by an FFT implementation in floating-point arithmetic with unit roundoff $u$.
The FFT is said to be normwise backward stable if there exists a constant $k_{N}>0$ such that, whenever $k_{N}u<1$, one can represent
$\widehat{y}=\mathcal{F}_{N}(x+\Delta x)\qquad\text{with}\qquad\|\Delta x\|_{2}\leq\gamma_{k_{N}}\,\|x\|_{2},\qquad\gamma_{k_{N}}:=\frac{k_{N}u}{1-k_{N}u}.$
Since $\mathcal{F}_{N}$ is unitary, the same bound holds in forward form:
$\|\widehat{y}-y\|_{2}\leq\gamma_{k_{N}}\,\|x\|_{2}.$
For the radix-$2$ Cooley–Tukey FFT ($N=2^{n}$), one may take $k_{N}=O(\log_{2}N)$; more precisely, if the twiddle factors are precomputed with absolute error bounded by $c_{N}u$, then a valid choice is
$k_{N}=\bigl(\sqrt{2}\,c_{N}+4+\sqrt{2}\bigr)\log_{2}N,$
see, e.g., [11].
Corollary 8.19 (Coefficient-wise enclosure from an $L^{2}$ bound).
Let $x\in\mathbb{C}^{N}$ be the vector of sampled values of a 22–periodic function $f$ on the uniform grid, let $\mathcal{F}_{N}(f)\in\mathbb{C}^{N}$ denote the exact discrete Fourier transform in the unitary normalization above, and let $\widetilde{\mathcal{F}}_{N}(f)$ be the output returned by FFTW after applying the corresponding $N^{-1/2}$ scaling. If $\|\widetilde{\mathcal{F}}_{N}(f)-\mathcal{F}_{N}(f)\|_{2}\leq\varepsilon_{2}$, then for each mode $\ell$,
$\bigl|\widetilde{\mathcal{F}}_{N}(f)[\ell]-\mathcal{F}_{N}(f)[\ell]\bigr|\leq\|\widetilde{\mathcal{F}}_{N}(f)-\mathcal{F}_{N}(f)\|_{\infty}\leq\|\widetilde{\mathcal{F}}_{N}(f)-\mathcal{F}_{N}(f)\|_{2}\leq\varepsilon_{2},$
and hence
$\mathcal{F}_{N}(f)[\ell]\in\bigl[\widetilde{\mathcal{F}}_{N}(f)[\ell]-\varepsilon_{2},\ \widetilde{\mathcal{F}}_{N}(f)[\ell]+\varepsilon_{2}\bigr].$
In particular, under the assumptions of Lemma 8.18 one may take $\varepsilon_{2}:=\gamma_{k_{N}}\,\|x\|_{2}$ (with $\gamma_{k_{N}}=k_{N}u/(1-k_{N}u)$).
Lemma 8.20.
Fix $k\in\mathbb{Z}$ and set $f_{k}:=e_{k}\circ T$. Let
$P_{k,\ell}:=\mathcal{F}(f_{k})[\ell]=\frac{1}{2}\int_{-1}^{1}f_{k}(x)\,e^{-\pi i\ell x}\,dx$
be the exact Fourier coefficient (in our normalization), and let $N\in\mathbb{N}$. Define the exact discrete coefficient
$\mathcal{F}_{N}(f_{k})[\ell]:=\frac{1}{N}\sum_{m=0}^{N-1}f_{k}(x_{m})\,e^{-\pi i\ell x_{m}},\qquad x_{m}:=-1+\frac{2m}{N}.$
Let $\widetilde{\mathcal{F}}_{N}(f_{k})[\ell]$ be the coefficient returned by FFTW.
Assume:
  1. (1)
    There exists $C_{k}>0$ such that $|\mathcal{F}(f_{k})[n]|\leq C_{k}/|n|^{2}$ for all $n\neq 0$.
  2. (2)
    Using interval arithmetic we compute enclosures $X_{m}\ni f_{k}(x_{m})$, and write $x_{m}^{\mathrm{mid}}:=\mathrm{mid}(X_{m})$ and $r_{m}:=\mathrm{rad}(X_{m})$. Let $x^{\mathrm{mid}}\in\mathbb{C}^{N}$ and $r\in\mathbb{R}_{+}^{N}$ be the corresponding vectors.
  3. (3)
    The FFT output satisfies $\|\widetilde{\mathcal{F}}_{N}(x^{\mathrm{mid}})-\mathcal{F}_{N}(x^{\mathrm{mid}})\|_{2}\leq\varepsilon_{2}$ for an explicitly computed $\varepsilon_{2}$.
Then, for every retained mode $|\ell|\leq N/2$,
(20) $\bigl|P_{k,\ell}-\widetilde{\mathcal{F}}_{N}(f_{k})[\ell]\bigr|\;\leq\;\underbrace{\frac{\pi^{2}}{3}\frac{C_{k}}{N^{2}}}_{\text{aliasing}}\;+\;\underbrace{\frac{\|r\|_{2}}{\sqrt{N}}}_{\text{interval sampling}}\;+\;\underbrace{\varepsilon_{2}}_{\text{FFT roundoff}}.$
Proof.
Decompose
$|P_{k,\ell}-\widetilde{\mathcal{F}}_{N}(f_{k})[\ell]|\leq$ $|P_{k,\ell}-\mathcal{F}_{N}(f_{k})[\ell]|+|\mathcal{F}_{N}(f_{k})[\ell]-\mathcal{F}_{N}(x^{\mathrm{mid}})[\ell]|$
$+|\mathcal{F}_{N}(x^{\mathrm{mid}})[\ell]-\widetilde{\mathcal{F}}_{N}(x^{\mathrm{mid}})[\ell]|.$
For 22–periodic $f_{k}$ one has the exact aliasing identity $\mathcal{F}_{N}(f_{k})[\ell]=\sum_{q\in\mathbb{Z}}\mathcal{F}(f_{k})[\ell+qN]$, hence
$|P_{k,\ell}-\mathcal{F}_{N}(f_{k})[\ell]|\leq\sum_{q\neq 0}|\mathcal{F}(f_{k})[\ell+qN]|\leq 2\sum_{q\geq 1}\frac{C_{k}}{(qN)^{2}}=\frac{\pi^{2}}{3}\frac{C_{k}}{N^{2}}.$
Let $\delta_{m}:=f_{k}(x_{m})-x_{m}^{\mathrm{mid}}$. Then $|\delta_{m}|\leq r_{m}$ and
$|\mathcal{F}_{N}(f_{k})[\ell]-\mathcal{F}_{N}(x^{\mathrm{mid}})[\ell]|=\Bigl|\frac{1}{N}\sum_{m=0}^{N-1}\delta_{m}\,e^{-\pi i\ell x_{m}}\Bigr|\leq\frac{1}{N}\sum_{m=0}^{N-1}|\delta_{m}|\leq\frac{\|\delta\|_{2}}{\sqrt{N}}\leq\frac{\|r\|_{2}}{\sqrt{N}}.$
By assumption (3) and $\|\cdot\|_{\infty}\leq\|\cdot\|_{2}$,
$|\mathcal{F}_{N}(x^{\mathrm{mid}})[\ell]-\widetilde{\mathcal{F}}_{N}(x^{\mathrm{mid}})[\ell]|\leq\|\mathcal{F}_{N}(x^{\mathrm{mid}})-\widetilde{\mathcal{F}}_{N}(x^{\mathrm{mid}})\|_{\infty}\leq\varepsilon_{2}.$
Combining the three estimates yields (20). ∎
Remark 8.21.
It is worth noting that, throughout, we take $N=2^{n}$ and rely on FFTW to compute the discrete Fourier transform. Our error bounds assume that FFTW uses twiddle factors computed with a precision comparable to machine precision [32]; accordingly, we include an a priori overestimation of $4u$, where $u$ denotes the unit roundoff. We also assume that the machine on which the experiments were run conforms to the IEEE 754 standard for floating-point arithmetic.
In principle, one could reimplement a radix-$2$ Cooley–Tukey FFT within our codebase (and even aim at a fully verified implementation), but absent a full formal development it is unlikely that such an implementation would match the performance, engineering maturity, and empirical validation of the FFTW development team. For this reason, we treat FFTW as a numerical kernel and compensate by incorporating explicit roundoff bounds (Lemma 8.18 and Corollary 8.19); see also the publicly documented FFTW/benchFFT accuracy benchmarking methodology and discussion [18, 17]. A further step towards end-to-end formal correctness would be to rely on formally generated code, for instance via verified LLVM generation from Isabelle/HOL as in [35].
@@ 9.1. Norm Bounds for the Discretized Operator
Recall that
$P_{\sigma,K}:=\pi_{K}N_{\sigma}P\pi_{K}$ and define the $L^{1}$–tail constant
$\|(1-\pi_{K})N_{\sigma}\|_{L^{1}\to L^{1}}\leq\Gamma^{(1)}_{\sigma,K}$
(cf. Lemma 8.7).
@@
For all $i\in\mathbb{N}$,
$\|P_{\sigma,K}^{i}\|_{L^{1}\to L^{1}}\leq(1+\Gamma^{(1)}_{\sigma,K})^{i}\,(2K+1).$
@@
Using $\pi_{K}^{2}=\pi_{K}$ (Lemma 8.3),
$P_{\sigma,K}^{i}=(\pi_{K}N_{\sigma}P)^{i}\,\pi_{K}.$
By Corollary 8.8 and $\|P\|_{L^{1}\to L^{1}}=1$,
$\|\pi_{K}N_{\sigma}P\|_{L^{1}\to L^{1}}\leq 1+\Gamma^{(1)}_{\sigma,K}$.
Moreover, with $\mathcal{F}(g)[j]=\frac{1}{2}\int_{-1}^{1}g(x)e^{-j\pi ix}\,dx$ we have $|\mathcal{F}(g)[j]|\leq\frac{1}{2}\|g\|_{L^{1}}$ and $\|e^{j\pi ix}\|_{L^{1}([-1,1])}=2$, hence
$\|\pi_{K}g\|_{L^{1}}\leq\sum_{j=-K}^{K}|\mathcal{F}(g)[j]|\;\|e^{j\pi ix}\|_{L^{1}}\leq(2K+1)\|g\|_{L^{1}}.$
Combining the estimates yields the claim. ∎
@@
For all $i\in\mathbb{N}$,
$\|P_{\sigma,K}^{i}\|_{L^{2}\to L^{2}}\leq\sqrt{2}\,\|N_{\sigma}\|_{L^{1}\to L^{2}}\,(1+\Gamma^{(1)}_{\sigma,K})^{i-1},$
where
$\|N_{\sigma}\|_{L^{1}\to L^{2}}=\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}\leq\sqrt{\frac{1}{2\sigma\sqrt{\pi}}\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr)}.$
Proof.
Set $A:=\pi_{K}N_{\sigma}P$. Using $\pi_{K}^{2}=\pi_{K}$, we have $P_{\sigma,K}^{i}=A^{i}\pi_{K}$. For $g\in L^{2}([-1,1])$, Cauchy–Schwarz and $\|\pi_{K}\|_{L^{2}\to L^{2}}\leq 1$ give
$\|\pi_{K}g\|_{L^{1}}\leq\sqrt{2}\,\|\pi_{K}g\|_{L^{2}}\leq\sqrt{2}\,\|g\|_{L^{2}}.$
Moreover, for $f\in L^{1}$,
$\|Af\|_{L^{1}}\leq\|\pi_{K}N_{\sigma}\|_{L^{1}\to L^{1}}\,\|P\|_{L^{1}\to L^{1}}\,\|f\|_{L^{1}}\leq(1+\Gamma^{(1)}_{\sigma,K})\|f\|_{L^{1}},$
and
$\|Af\|_{L^{2}}\leq\|\pi_{K}\|_{L^{2}\to L^{2}}\,\|N_{\sigma}\|_{L^{1}\to L^{2}}\,\|P\|_{L^{1}\to L^{1}}\,\|f\|_{L^{1}}\leq\|N_{\sigma}\|_{L^{1}\to L^{2}}\,\|f\|_{L^{1}}.$
Therefore,
$\|P_{\sigma,K}^{i}g\|_{L^{2}}=\|A^{i}\pi_{K}g\|_{L^{2}}\leq\sqrt{2}\,\|N_{\sigma}\|_{L^{1}\to L^{2}}\,(1+\Gamma^{(1)}_{\sigma,K})^{i-1}\,\|g\|_{L^{2}},$
which proves the operator bound. The expression for $\|N_{\sigma}\|_{L^{1}\to L^{2}}$ is Young’s inequality, and the coth bound follows from Lemma 8.17 and Remark 8.16. ∎
@@
There exist $K,N\in\mathbb{N}$ such that for all $k>K$,
$\|P_{\sigma,k}^{N}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq\frac{3}{4}.$
Proof.
Choose NN so that $\|P_{\sigma}^{N}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq\frac{1}{2}$ (Proposition 6.7). For $g\in\mathcal{U}_{0}$,
$\|P_{\sigma,k}^{N}g\|_{L^{2}}\leq\|P_{\sigma}^{N}g\|_{L^{2}}+\|(P_{\sigma}^{N}-P_{\sigma,k}^{N})g\|_{L^{2}}\leq\frac{1}{2}\|g\|_{L^{2}}+\|P_{\sigma}^{N}-P_{\sigma,k}^{N}\|_{L^{2}\to L^{2}}\,\|g\|_{L^{2}}.$
Now,
[…]
We can now bound each term using Theorem 8.4 together with the $L^{1}$ growth bound in Lemma 9.1. The resulting discretization error is controlled by the Fourier tail of the Gaussian multiplier and decays exponentially in $k$, so for $k$ large enough one gets $\|P_{\sigma}^{N}-P_{\sigma,k}^{N}\|_{L^{2}\to L^{2}}\leq\frac{1}{4}$, proving the claim. ∎
@@
Let $f_{\sigma}$ be the fixed point of $P_{\sigma}$, and $f_{\sigma,K}$ the fixed point of $P_{\sigma,K}$. Then for any $\epsilon>0$ there exists $K$ such that
@@
Apply Theorem 8.12 to the pair of operators $P_{\sigma}$ and $P_{\sigma,K}$, using the uniform contraction on $\mathcal{U}_{0}$ provided by Lemma 9.3. The perturbation size is measured by the discretization error
$\|P_{\sigma}-P_{\sigma,K}\|_{L^{1}\to L^{2}}\;\;\leq\;\;\|(1-\pi_{K})N_{\sigma}\|_{L^{1}\to L^{2}}\;+\;\|\pi_{K}(N_{\sigma}P-\!N_{\sigma}P\pi_{K})\|_{L^{1}\to L^{2}},$
which is controlled by the Fourier tail of the Gaussian multiplier and decays exponentially in $K$ Lemma 8.4. The remaining operator factors in Theorem 8.12 are bounded by Lemma 9.2, which gives
$\|P_{\sigma,K}^{i}\|_{L^{2}\to L^{2}}\leq\sqrt{2}\,\|N_{\sigma}\|_{L^{1}\to L^{2}}\,(1+\Gamma^{(1)}_{\sigma,K})^{i-1},\qquad\|N_{\sigma}\|_{L^{1}\to L^{2}}=\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}$
note that Lemma 9.2 yields the $C_{i}$ for Theorem 8.12. Since $\Gamma^{(1)}_{\sigma,K}\to 0$ exponentially as $K\to\infty$, the error bound furnished by Theorem 8.12 can be made smaller than $\epsilon$ by taking $K$ sufficiently large, proving the claim. ∎
@@
Let $\phi\in L^{2}(m)$. Then the Birkhoff average
@@
converges almost surely to $\int\phi\,f_{\sigma}\,dm$. Moreover, for any $\epsilon>0$ there exists $K$ such that
$\Bigl|\int\phi\,f_{\sigma}\,dm-\int\phi\,f_{\sigma,K}\,dm\Bigr|<\epsilon.$
@@
Uniqueness of the stationary measure (Proposition 6.8) yields ergodicity, hence the almost-sure convergence follows from the Birkhoff ergodic theorem (Theorem 5.13). For the approximation error, by Cauchy–Schwarz,
$\Bigl|\int\phi\,(f_{\sigma}-f_{\sigma,K})\,dm\Bigr|\leq\|\phi\|_{L^{2}}\,\|f_{\sigma}-f_{\sigma,K}\|_{L^{2}},$
and the claim follows from Theorem 9.4. ∎